A certain population is known to be growing at a rate given by the logistic equation

dx/dt =x(b - ax)

Show that the minimum rate of growth will occur when the population is equal to half the

equilibrium size, that is, when the population is b / 2a .


1
Expert's answer
2021-05-28T10:15:29-0400

Let's find the minimum rate of growth, so the function "x(b-ax)" should be minimized. The


derivative of the rate change must be zero:


"\dfrac{d(bx-ax^2)}{dx} = 0"


"b-2ax = 0"


"x = \dfrac{b}{2a}"


"\dfrac{d^2(bx-ax^2)}{dx^2} = -2a<0" if "a>0" , hence in this case we get a maximum;


"\dfrac{d^2(bx-ax^2)}{dx^2} = -2a>0" if "a<0" , hence in this case we get a minimum.


Such a population x is equal to half the equilibrium size.